OpenStudy (anonymous):

A man who is heterozygous for tongue rolling and has a normal vision and a woman who cannot roll her tongue and is heterozygous for colorblindness have a child. What is the probability that it will be able to roll his/her tongue and have normal vision? Tongue-rolling is due to a dominant allele. Color Blindness is caused by an X-linked recessive gene.

OpenStudy (anonymous):

Tongue-rolling is due to a dominant allele- phenotype of father-Rr phenotype of mother-rr phenotype of children-Rr,Rr,rr,rr 50% chance of tongue roll in children. colour blindness x linked recessive--->25% chance of a child with colour blindness. The probability that it will be able to roll his/her tongue and have normal vision might be 25% ----Not sure .........................sorry Forgot every thing........

OpenStudy (anonymous):

thank you for taking your time to answer! I know many people hate genetics.. haha

OpenStudy (anonymous):

The probability that a child born to these parents is able to roll their tongue and have normal vision is 1/8. Probability of tongue roll is 1/2 (or 2/4) and probability of color blindness is 1/4. So you multiply these probabilities to find the probability that the child has both. 2/4 * 1/4 = 2/16 or 1/8

OpenStudy (anonymous):

men( dominant) female (recessive ) X(R)-Y X(r)-X(r) = 2 X(R) X(r) +2Y X(r) so the girls will not have the blind and all the sons will have the blind that's what i think because the dominant R from the father will stop the recessive one in girls but the Y from the father as we know doesn't not have an allele .. so the allele will only come from the mom in males !